if the length of a rectangle is 4 units more
than its breadth
if its perimetre is 64
units find its length and breadth???

Answer by mananth(16949)

if the length of a rectangle is 4 units more
than its breadth
if its perimetre is 64
units find its length and breadth???
...
breadth = x
length = x+4
2*(L+B)= perimeter
2*(x+x+4)=64
2(2x+4)=64
/2
2x+4 =32
2x=32-4
2x=28
x=14 the breadth
length = 14 +4 = 18
...
m.ananth@hotmail.ca